What is significant about the difference between the domain of y= square root of x and y= x2?


1
Expert's answer
2021-03-17T17:36:04-0400

given that :

1). "y=\sqrt{x}"


domain = "[0,\infty)" and range = "[0,\infty)"


2). "y=x^2"


domain = "(-\infty,+\infty)" and range = "[0,\infty)"


both graph have different domain but have same range, this is because as in first graph "\sqrt{x}" is present which is taken only the positive values of x for obtaining the real values range of graph. Thus, the domain of this graph is "[0,\infty)" and gives values as range is "[0,\infty)" .


But in the second graph x2 is present and due to square nature it can take all real values either positive or negative as domain but its range is "[0,\infty)", as all the negative values of the x turn to positive value due to squaring so it doesn't gives negative values.


Function is given as:

"y=\sqrt[2]{x^{-2}+4}"


"y=\sqrt[2]{\frac{1}{x^2}+4}"


for real solutions of the term inside the square root should be as:

"\frac{1}{x^2}+4\ge0"


We can se that has degree two so for all values of x the above term gives always positive values except x=0, as this makes it infinite value i.e. undefined, so the domain of the given equation is as:

Domain = ("-\infty,+\infty" )-[0]

Function is given as:

"y=\sqrt[2]{x^{-2}+4}"


"y=\sqrt[2]{\frac{1}{x^2}+4}"


for real solutions of the term inside the square root should be as:

"\frac{1}{x^2}+4\ge0"


We can se that has degree two so for all values of x the above term gives always positive values except x=0, as this makes it infinite value i.e. undefined, so the domain of the given equation is as:

Domain = ("-\infty,+\infty" )-[0]